# Forecasting Response to Treatment with Global Deep Learning and Patient-Specific Pharmacokinetic Priors

### v6

@@ A.1 Data
Fig. 7 shows the first 10 days of data for 6 subjects from the simulated dataset.
Fig. 8 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset.
Table 2 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
Table 2: Sample count (miss- ing samples) for the original OhioT1DM datasets and prepro- cessed test set
@@
Figure 7: Time series examples for the first 10 days of 6 subjects from the simulated dataset.
@@ A.2 Models
Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
Neural Hierarchical Interpolation for Time Series (NHITS) - A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.(Elman, 1990; Cho et al., 2014)
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies. TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
A.3 Hyperparameters
For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 3. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 3.
@@ A.3 Hyperparameters
Learning rate Loguniform(1e-4, 1e-2)
@@
Dropout {0.0, 0.1, 0.2, 0.3, 0.4, 0.5}
Training steps {1000, 2000}
@@
Basal dose initial $k_{a}$ Uniform(1, 2, 11)
Bolus dose initial $k_{a}$ Uniform(1, 2, 11)
@@
We tune additional specific hyperparameters for each architecture. For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the TFT the hidden size ([128,256]); for the NHITS we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
@@ A.4 Metrics
Appendix B Second Appendix
B.1 Patient-specific (local) models
@@
We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects. Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table. 4 and Fig 9b. Furthermore, the global NHITS-Pharamcokinetic model consistently outperforms the global NHITS-Exogenous model, on average, for each OhioT1DM subject as shown in Fig 9a.
@@
Figure 9: The pharmacokinetic NHITS  model outperforms the NHITS baseline exogenous model across all patients, on average (a). Additionally, the pharmacokinetic model trained on data from subjects, or the global model, achieves lower MAE than models trained on individual subjects, or local models (b).
B.2 Forecasting treatment effects
In this appendix, we explore the effect of changing the insulin dosage on the forecasts produced by our approach. Fig. 10, demonstrates that the pharmacokinetic model appropriately learns the treatment effect of insulin on blood glucose for OhioT1DM subjects, decreasing the forecast when the dosage increases.
Figure 10: Treatment effect of insulin on the NHITS pharmacokinetic model. Each point in the plot corresponds to the final horizon point. Forecasts with original bolus doses (orange), removing bolus insulin doses results in higher forecasts (blue), and augmenting doses results in lower blood glucose forecasts (yellow).
B.3 Model MAE and RMSE results
This appendix presents the complete results for all models, metrics of evaluation, and both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis.

### v7

@@ A.1 Data
Simulated data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) [Xie, 2018, Visentin et al., 2016] provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes. Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. In a similar approach to [Rubin-Falcone et al., 2022], the meal schedule used to generate simulated data was based on the Harrison-Benedict equation [Arthur Harris and Gano Benedict, 21919] with approximately 3 meals per day and no additional snacks. Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
The ‘Dexcom’ continuous glucose monitor (CGM) option was used to obtain glucose readings at 5-minute intervals. The default basal-bolus controller provided in the simulation was used to administer insulin at the time of each meal, where the bolus amount is computed based on the current glucose level, the target glucose level, the patient’s correction factor, and the patient’s carbohydrate ratio. However, we adapt the controller to deliver basal doses once per hour, consistent with the function of real insulin pumps used in the OhioT1DM dataset. The resulting simulated data consists of CGM (in mg/dL) values, insulin (units), and carbohydrates (in grams), referred to as CHO. We extract the bolus insulin rate from the overall insulin rate as the difference in the continuous basal insulin rate, which is consistent hourly doses. Blood glucose time series for 6 simulated patients are shown in Fig. 7 in Appendix A.1. Basal and bolus dose distributions for 12 simuluted patients are shown in Fig. 9 in Appendix A.1.
OhioT1DM
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020]. Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
Sparse exogenous features in the data include basal rate (in units per hour), temporary basal rate (units per hour), and bolus insulin (units), and CHO values. Bolus insulin consists of both ‘normal’ insulin, which is recorded as an instantaneous administration, and ‘square dual’, where the insulin dose is stretched over a defined period. ‘Square dual’ insulin doses were divided evenly across the specified administration window with the assumption that insulin units are administered consistently across specified timestamps. If multiple bolus administrations occurred within the same minute, the insulin units were summed. For Medtronic insulin pumps 530G and 630G versions, basal insulin is administered hourly based on the basal rate. For timestamps with overlapping basal rates, the last basal rate value was considered. When specified, the temporary basal rate superseded the basal rate.
We use the specified train and test partitions provided in the dataset. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1. Missing values were forward filled to prevent data leakage. Blood glucose time series plots for 6 subjects are shown in Fig. 8 in Appendix A.1. Basal and bolus dose distributions for the 12 patients are shown in Fig. 10 in Appendix A.1.
Fig. 7 shows the first 10 days of data for 6 subjects from the simulated dataset. Fig. 8 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset. Figs. 9 and 10 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively. Table 2 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
Table 2: Sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set
@@ A.2 Models
The baseline models used in this work have been extensively used in previous blood glucose forecasting work. For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies [Qian et al., 2021, Li et al., 2019a, Fox et al., 2018, Rabby et al., 2021, Rubin-Falcone et al., 2020, 2022, Miller et al., 2020]. Similarly, CNNs [Xie and Wang, 2020, Li et al., 2019a, b] and MLPs [McShinsky and Marshall, 2020, Miller et al., 2020] have been utilized in proposed models and as baselines. NHITS achieves state-of-the-art performance, surpassing RNN-based, transformer-based, and other deep learning time series decomposition models (such as N-BEATS [Oreshkin et al., 2020]) in long-horizon forecasting tasks [Challu et al., 2022].
Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. [Hyndman, 2008].
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation [Nair and Hinton, 2010, Fukushima, 1975, Rosenblatt, 1958].
Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation [Challu et al., 2022].
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.[Elman, 1990, Cho et al., 2014]
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.[Sak et al., 2014]
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions [Bai et al., 2018, van den Oord et al., 2016].
A.3 Model Training and Hyperparameters
Models were trained using Adam optimizer [Kingma and Ba, 2017]. Optimal hyperparameters were selected based on a validation set with cross-validation. Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 3. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 3. We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
@@ A.3 Model Training and Hyperparameters
Learning rate Loguniform(1e-4, 1e-1)
@@
Windows batch size’ 256
Dropout 0.0
Training steps 2000
Validation check steps 100
Early stop 5
patience steps
@@
Initial $\textbf{k}_{\text{basal}}$1.0, 1.1, 1.2}
Initial $\textbf{k}_{\text{bolus}}$1.7, 1.8, 1.9}
Initial $\textbf{k}_{\text{CHO}}$1.7, 1.8, 1.9}
@@
We tune additional specific hyperparameters for each architecture. For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the NHITS we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
@@ A.4 Metrics
We use the following metrics to evaluate model performance. Here, $H$ refers to the forecast horizon.
@@
A.5 Propositions
Proposition A.1.
Under linear pharmacokinetic (PK) assumptions, the area under the concentration curve is directly proportional to the dose of the drug. As such, PK parameters, such as bioavailability, are assumed to be constant for a given drug. The bioavailability, $F$, can be assessed from the dose-normalized area under the concentration curve (AUC). When bioavailability is constant, we show that the AUC is bounded by the dose, $x$, given that $0\leq F\leq 1$,
$F$ $\propto\frac{\text{AUC}}{x}$ (9)
$n\cdot F$ $=n\cdot\frac{\text{AUC}}{x}$ (10)
AUC $=F\cdot x$ (11)
$\leq x$ (12)
where $n$ where is a constant of proportionality.
Proposition A.2.
Given some mathematical model, $C(t,\cdot)_{x}$, of concentration for a drug dose, $x$, that is a function of time, $t$, and pharmacokinetic parameters, $\cdot$, we show that the overall effect (concentration) of $n$ drug doses of the same drug is equal to the sum of the individual effects under linear pharmacokinetic assumptions:
[…]
$=\int_{0}^{T}\sum_{i=1}^{n}C(t,\cdot)_{x}dt$ (17)
Here, $T$ represents the total duration of drug exposure or the time interval over which drug concentrations are being measured and $t$ represents a specific point in time within that interval.
A.6 Algorithm
We outline the algorithm for training our proposed PK encoder that transforms sparse exogenous variables into time-dependent concentration profiles. This framework can be implemented in any deep learning architecture. It can also be adapted to utilize other concentration-time profile functions pertinent to a specific healthcare context.
The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch. This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library [Liaw et al., 2018]. For each training step, the algorithm iterates over batches of data. It computes a context vector c based on the input data x, k, and other parameters. It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values. After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step. It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$. It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss. Finally, the algorithm returns the final k values.
Algorithm 1 PK Encoder Training.
$E$ is the number of trainings steps, $\eta$ is the learning rate, $N$ is the number of subjects, $H$ is the forecast horizon, and $L$ is the lag time.
Data: $N>0,f\in\mathbb{R}$, $L\in\mathbb{R}$, $H\in\mathbb{R}$
Result: k
$\mathcal{B}_{train}\leftarrow$ (split training data into batches)
$\mathcal{B}_{val}\leftarrow$ (split validation data into batches)
$\textbf{k}^{*}\leftarrow$ (Randomly select $\textbf{k}^{*}$ from defined hyperparmeter options, $\textbf{k}^{*}\in\{1,\ldots,2\}$)
for each step i from 1 to E do
       $\textbf{k}\leftarrow\textbf{k}^{*}$triangleright Initialize k
for batch $b\in\mathcal{B}_{train}$ do
               $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
$\textbf{k}\leftarrow\textbf{k}-\eta\cdot\nabla_{\textbf{k}}\ell(\textbf{k})$triangleright Update k using SGD
       end for
      if i % validation step == 0 then
               $\ell(\textbf{k})\leftarrow 0$
for batch $b\in\mathcal{B}_{val}$ do
                     $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
             end for
            $\textbf{k}^{*}\leftarrow\arg\min_{\textbf{k}^{*}\in\{1,\ldots,2\}}\ell(\textbf{k})_{total}$triangleright Select the hyperparameter that minimizes loss
       end if
end for
return k
Figure 7: Time series for the first 10 days of 6 subjects from the simulated dataset.
Figure 9: Density plots of bolus and basal insulin for 12 simulated data subjects. Bolus doses are general larger than basal doses.
Figure 10: Density plots of bolus and basal insulin for OhioT1DM subjects. Bolus doses are general larger than basal doses.
Appendix B Pharmacokinetic Background
Concentration-time Profiles
From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects [Rosenbaum, 2017, Angelika et al., 2005]. Alternatively, subcutaneous injections involve absorption and elimination processes where the peak onset time may not be instantaneous depending on the drug profile. Regarding knowledge on concentration-time profiles established through empirical and invasive studies, we know that concentration-time profiles for subcutaneous injections adhere to right-skewed curves. The degree of right-skew in the concentration curve will differ depending on the type of insulin, as illustrated in Fig. 2, along with various patient-specific factors.
Insulin Analogs
Insulin analogs are modified forms of insulin designed to exhibit certain PK properties, such as changes in drug onset and duration of action. Fig. 2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs [Angelika et al., 2005]. Generally, bolus doses are quick-acting and, thus, often taken before meals. Contrastingly, basal insulin is longer-acting and serves to regulate blood glucose levels throughout the day. However, continuous glucose monitors (CGMs) and insulin pumps generally only hold and administer one type of insulin, such as rapid-acting insulin. For these devices, ‘bolus’ and ‘basal’ are terms used to refer to different timings of injections for different purposes. Bolus doses are administered discretely, generally around meals, while basal doses are administered continuously at a defined frequency, such as hourly, to regulate blood glucose levels throughout the day.
Bioavailability
In Fig. 2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Bioavailability varies based on many factors, including type of medication, dosage, and patient-specific characteristics. For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent [Søeborg et al., 2012]. However, for subcutaneous injections, bioavailability is typically lower due to unintended drug absorption or elimination.
First-order Kinetics
First-order kinetics, a common assumption in pharmacokinetics, describes the rate at which a drug is eliminated from the body as being proportional to its concentration. This means that a constant fraction of the drug is eliminated per unit of time, regardless of its concentration levels. Exponential functions are often used to represent concentration-time profiles under the assumption of first-order kinetics, simplifying the modeling of plasma drug concentration in the body.
Linear Pharmacokinetics
Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear [Rosenbaum, 2017]. In linear pharmacokinetics, the area under the concentration curve is directly proportional to the dose of the drug. If you double the dose of a drug, the area under the curve will also double. Linearity results from the assumption that the drug processes exhibit first-order kinetics and that the PK parameters remain constant with dose, indicating no significant changes in the body’s ability to eliminate the drug with higher doses. The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics [Rosenbaum, 2017].
Figure 11: When doses of insulin are administered at close intervals, the subsequent dose may start to act before the previous dose has fully finished its effect. This can lead to a stacking of insulin effects. The figure depicts additive effects under linear pharmacokinetic assumptions.
Appendix C Additional Results
C.1 Patient-specific (Local) Models
@@
We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects. Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table. 4 and Fig. 5. Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig. 12.
@@
All 10.324 8.873 10.425 8.758
(0.401) (0.064) (0.659) (0.080)
Figure 12: The NHITS-PK hybrid global-local model (Hybrid) trained on all OhioT1DM subject data (blue, solid) has a lower MAE, on average, than NHITS-PK models trained on individual subjects, or local models (purple, dashed) .
C.2 Forecast Comparison
Figure 13: Blood glucose forecasts of the final horizon point during 5 randomly selected hypoglycemic threshold periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients. The NHITS-PK hybrid global-local model (orange, dashed) produces improved forecasts of the ground truth (blue, solid) compared to the NHITS baseline model with sparse exogenous features (yellow, dash-dotted).
C.3 Forecasting Response to Treatment
Predicting treatment effects using accurate forecasting algorithms can assist individuals with diabetes and their care providers in making informed decisions about the ideal medication type and dosage, leading to improved glycemic control and reduced risk of both short- and long-term complications. Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts. We examine whether the NHITS-PK model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10. When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig. 14. The NHITS-PK model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
Figure 14: Treatment effect of insulin on the NHITS-PK hybrid global-local model. Each point in the plot corresponds to the final horizon point. Forecasts with original bolus doses (orange, solid), removing bolus insulin doses results in higher forecasts (blue, dashed), and augmenting doses results in lower blood glucose forecasts (yellow, dotted).
C.4 Forecasting with Dose-Independent Parameters
[…]
We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values. However, this approach led to worse performance on the OhioT1DM dataset. The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters [Xie, 2018, Visentin et al., 2016]. In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3. Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data. These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
Table 5: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
Simulated Dataset
Model All Values Critical Values
Baseline Exog. Baseline Exog.
NHITS 8.268 7.304 11.114 9.094
(0.039) (0.055) (0.237) (0.236)
NHITS-ST 7.055 8.648
(0.038) (0.068)
NHITS-PK 7.103 8.449
linear (0.069) (0.139)
NHITS-PK 7.037 8.492
nonlinear (0.093) (0.220)
OhioT1DM Dataset
NHITS 9.060 8.873 10.337 10.137
(0.142) (0.064) (0.137) (0.108)
NHITS-ST 8.958 10.131
(0.087 ) (0.100)
NHITS-PK 8.830 10.026
linear (0.060) (0.064)
NHITS-PK 8.758 9.965
nonlinear (0.080) (0.089)
C.5 Model MAE and RMSE Results
This appendix presents the complete results for all models for both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis. Table 6 provides complete results for the simulated data and Table 7 provides complete results for the OhioT1DM dataset.